# Permutation-based multiple testing when fitting many generalized linear models

### v1

@@ Abstract
The multiple testing problem appears when fitting multivariate generalized linear models for high dimensional data. We show that the sign-flip test can be combined with permutation-based procedures for assessing the multiple testing problem
@@ 1 Introduction
In high-dimensional data, such as neuroimaging and transcriptomics, it is common to fit many generalized linear regression models in parallel, each with a small sample size (Schaarschmidt et al., 2022; Love et al., 2014; Winkler et al., 2014). Usually, the goal of the analysis is to perform hypothesis testing to find relevant associations, usually after adjusting the p-values for multiple testing.
@@
This approach comes with several challenges. In the first place, small sample sizes can make classical tests unreliable, especially in nonlinear models where exact methods are not available and tests rely on asymptotic arguments. In such cases, the usual normal approximation of the test statistic can be quite unreliable (Schaarschmidt et al., 2022). Secondly, the generalized linear model makes some crucial assumptions which are difficult to check, especially for nonlinear models with small sample size. In particular, the detection of overdispersion over the assumed models can be quite problematic, especially since the variance is not constant among the observations even without overdispersion. Proper model checking is further hampered by the sheer number of models that need to be checked in high-dimensional data. Finally, the test statistics of the parallel models are often correlated, due to correlations in the underlying biological measurements. Classical multiple testing corrections (such as Bonferroni-Holm) are designed in order to protect against any correlation structure, but can be very conservative in the presence of such correlations (Goeman and Solari, 2014; Gao et al., 2010; Saffari et al., 2018), while taking correlations into account could increase the power of the test procedure.
@@
In order to address all these issues we propose a permutation-based multiple testing procedure in combination with the sign-flip score test (Hemerik et al., 2020; De Santis et al., 2022). The test shows reliable behavior for small sample sizes, and has been shown to be robust against general variance misspecification under minimal assumptions. We start by defining a test for a global null hypothesis about a multivariate regression parameter, which guarantees weak control of the familywise error rate (FWER). The global test statistics can be used to make additional inference through the Closed Testing approach (Marcus et al., 1976). Consequently we compute adjusted p-values for each individual hypothesis through a multiple testing procedure based on the max-$T$ method of Westfall and Young (1993), which guarantees strong control of the FWER, that is, it produces valid adjusted p-values for all possible subsetting hypotheses.
The key point of using a permutation-based solution is that the proposed method adapts to the unknown correlation structure. It is especially useful when strong correlation between the individual test statistics is present; in that situation it guarantees a relevant gain in power over alternative methods, as we will show in a simulation. Further, the sign-flip test can be applied to estimate a lower bound for the true discovery proportion (TDP), that is, the proportion of non-null coefficients over a pre-specified group of hypotheses, using a permutation-based framework (Goeman and Solari, 2011; Andreella et al., 2023; Vesely et al., 2023; Blain et al., 2022).
@@
The paper is organized as follows: Section 2 revisits the sign-flip score test for univariate testing; Section 3 contains the novel contribution of the paper, that is, the multivariate testing extension. Section 4 contains a simulation study.
@@
Code to implement our methods is available in the flipscores R package (R Core Team, 2023), available from CRAN and in the pyperm python package (Davenport, 2023). Code to reproduce the results of this paper is available at github.com/rds/multiflip.
@@ 2 One dependent variable
In this Section we review the derivation of the univariate sign-flip score test as detailed in Hemerik et al. (2020) and De Santis et al. (2022).
Let $Y_{i}$ be the target variable which belongs to the exponential dispersion family, i.e., with a density of the form (Agresti, 2015)
$f_{\beta,{\bm{\gamma}},\bm{X}_{i}}(y_{i})=\exp\left\{\frac{y_{i}\theta_{i}-b(\theta_{i})}{a(\phi_{i})}+c(y_{i},\phi_{i})\right\},$
@@
where $g(\cdot)$ is the link function, the covariate $X$ is an $n\times 1$ matrix, i.e., a column vector, such that $\beta\in\mathbb{R}$. Further, the nuisance covariates $Z$ are a $n\times(k-1)$ matrix.
Throughout the section we are interested in the null hypothesis $H_{0}:\beta=\beta_{0}$ against a one or two-sided alternatives. Consequently, the other parameters $\gamma$ (and $\phi_{i})$ are considered as nuisance parameters.
Let
@@
The diagonal matrix of the GLM weights is defined as $W=DV^{-1}D$ with entries $w_{i}=({\partial\mu_{i}}/{\partial\eta_{i}})^{2}/var(Y_{i})$ (Agresti, 2015). Finally,
$H={W}^{1/2}Z(Z^{\prime}{W}Z)^{-1}Z^{\prime}{W}^{1/2}$
represents the projection matrix for GLMs.
The test is based on the effective score. The interpretation of the effective score is that if the true nuisance parameter is plugged in, the effective score is the residual from the projection of the marginal score for $\beta$ on the space spanned by the nuisance scores (Marohn, 2002). If the estimated nuisance parameter $\hat{\gamma}$ is plugged in, the effective score is
$n^{1/2}S^{*}_{\hat{\gamma}}=X^{\prime}{W}^{1/2}(I-{H}){V}^{-1/2}(Y-\hat{\mu})=\sum_{i=1}^{n}\nu^{*}_{\hat{\gamma},i}.$
Note that indeed the effective score can be written as a sum of $n$ elements $\nu^{*}_{\hat{\gamma},i}$, which we call the score contributions.
Regarding the estimator $\hat{\bm{\gamma}}$, we make the same assumptions of De Santis et al. (2022), in particular that $\hat{\bm{\gamma}}$ is a $\sqrt{n}$-consistent estimate of the true parameter $\bm{\gamma}_{0}$.
The test is performed by means of random sign flipping of score contributions. Define $g_{1}=(1,\dots,1)\in\mathbb{R}^{n}$ and for every $2\leq j\leq w$ let $g_{j}=(g_{j1},\dots,g_{jn})$ be independent and uniformly distributed on $\{-1,1\}^{n}$. For $1\leq j\leq w$, let the superscript $j$ denote that $g_{j}$ has been applied. The transformed test statistics are defined as
$S^{*j}_{\hat{\bm{\gamma}}}$ $=n^{-1/2}\sum_{i=1}^{n}g_{ji}\nu^{*}_{\hat{\bm{\gamma}},i}.$
The derivation of the flipped effective score statistic in this matrix notation is straightforward. Note that the effective score is the product of an $1\times n$ and an $n\times 1$ vector. This means that we can write it multiplying the $n\times 1$ matrix by the $n\times n$ sign-flipping matrix $G_{j}$, with diagonal entries $g_{j1},\dots,g_{jn}$. The transformed test statistic is therefore
$S^{*,j}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}{W}^{1/2}(I-{H}){V}^{-1/2}G_{j}(Y-\hat{\mu}).$
@@
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}=S^{*j}_{\hat{\bm{\gamma}}}$ and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{0}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
We will now recall the definition of the test proposed in De Santis et al. (2022), which is a recent upgrade of the test above. This recent adaptation often improves the small-sample performance of the test by standardization of the effective score. We call the resulting test statistic the standardized score statistic, defined as
$S^{S}_{\hat{\bm{\gamma}}}=S^{*}_{\hat{\bm{\gamma}}}/Var\{S^{*}_{\hat{\bm{\gamma}}}\}^{1/2},$ (1)
@@
$Var\{S^{*}_{\hat{\bm{\gamma}}}\}=n^{-1}X^{T}W^{1/2}(I-H)W^{1/2}X.$
The transformed test statistic, for a generic flip $G_{j}$, is defined as
$S^{S,j}_{\hat{\bm{\gamma}}}=S^{*,j}_{\hat{\bm{\gamma}}}/Var\{S^{*,j}_{\hat{\bm{\gamma}}}\}^{1/2},$ (2)
@@
$Var\{S^{*,j}_{\hat{\bm{\gamma}}}\}=n^{-1}X^{T}W^{1/2}(I-H)G_{j}(I-H)G_{j}(I-H)W^{1/2}X.$
@@
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}=S^{S,j}_{\hat{\bm{\gamma}}}$ and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{0}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
Some robustness properties of this test are discussed in Hemerik et al. (2020) and in De Santis et al. (2022). In particular, the proposed test is proven to be asymptotically exact in case of general variance misspecification under minimal assumptions.
3 Multiple responses
Suppose there are $m\geq 2$ dependent variables $Y^{1},\dots,Y^{m}$. For each response $Y^{l}$ we have $n$ independent observations $Y^{l}_{1},\dots,Y^{l}_{n}$, which follow some model in the exponential dispersion family. We consider $m$ null hypotheses $H_{1},\dots,H_{m}$, where $H_{l}$ is the hypothesis that $\beta^{l}=\beta^{l}_{0}$. Like before, we assume $\hat{\bm{\gamma}}^{l}$ is a $\sqrt{n}$-consistent estimate of $\bm{\gamma}_{0}^{l}$.
@@ 3 Multiple responses
$S^{*,l}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}({W}^{l})^{1/2}(I-{H}^{l})({V}^{l})^{-1/2}(Y^{l}-\hat{\mu}^{l})=n^{-1/2}\sum_{i=1}^{n}\nu^{*l}_{\hat{\bm{\gamma}},i}.$
Analogously to Section 2, we define the sign-flipped effective score test statistic for a generic flip matrix $G_{j}$ as
$S^{*,j,l}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}({W}^{l})^{1/2}(I-{H}^{l})({V}^{l})^{-1/2}G_{j}(Y^{l}-\hat{\mu}^{l})=n^{-1/2}\sum_{i=1}^{n}g_{ji}\nu^{*l}_{\hat{\bm{\gamma}},i}.$
Further, we have
$Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}=X^{\prime}({W}^{l})^{1/2}(I-{H^{l}})G_{j}(I-{H^{l}})G_{j}(I-{H^{l}})({W}^{l})^{1/2}X.$
while the standardized score is
$S^{S,j,l}_{\hat{\bm{\gamma}}}=S^{*,j,l}_{\hat{\bm{\gamma}}}Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}^{-1/2}.$
@@
$H_{0}=\bigcap_{l\in L}H_{L}:\beta^{l}=\beta^{l}_{0}$ (3)
where $L$ is a set containing any combination of the $m$ hypotheses. Note that $H_{0}$ can be either a global or a partial null hypothesis for a subset of parameters. For simplicity, we will derive the procedure for the global null hypothesis, but it is exactly analogous for partial null hypotheses of any kind.
A first idea to perform the test might be given by a full standardization. We can indeed build a global test statistic based on the form of a Mahalanobis distance as follows. Let $S^{*,j,L}_{\hat{\bm{\gamma}}}$ be the vector with element $S^{*,j,l}_{\hat{\bm{\gamma}}}$, $1\leq l\leq m$. Note that by definition it is an $m$-dimensional vector. Call $Var\{S^{*,j,L}_{\hat{\bm{\gamma}}}\}$ the corresponding variance-covariance matrix, which has dimension $m\times m$ and is assumed (for the moment) known. Let the joint test statistic be
$T_{j,L}^{n}=\left(S^{*,j,L}_{\hat{\bm{\gamma}}}\right)^{\prime}Var\left\{S^{*,j,L}_{\hat{\bm{\gamma}}}\right\}^{-1}\left(S^{*,j,L}_{\hat{\bm{\gamma}}}\right)$ (4)
If we assume the normality of the response we have the following theorem;
Theorem 3.
Assume to fit a normal linear regression model. Let $T_{j,L}^{n},1\leq j\leq w$ as defined in (4). Let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. Under $H_{0}$, the test is an exact $\alpha$ level test.
Proof.
Trivially, note that $S^{*,j,L}_{\hat{\bm{\gamma}}}$ is normally distributed with zero mean and that the test statistic is a quadratic form. The expected value of each test statistic is equal to $m$ for every flip, while the variance is equal to $2m$ for every flip. This follows from standard properties of random vectors. It implies that all the test statistics share the first two moments. Thus we can directly apply Theorem 1 of Hemerik and Goeman (2018) to show that the test derived is an $\alpha$-level test for finite sample size. ∎
The procedure outlined has some issues. First of all, the test is exact only for normal responses with known variance. Otherwise, we have to follow asymptotic arguments related to the asymptotic normal distribution of the effective score statistic (Marohn, 2002). Another weak point is the fact that for each flip it is required to invert an $m\times m$ matrix for each flip, which might be unfeasible for large values of $m$. Further, we have to estimate the correlation between responses, which must be assumed to be known except for a limited number of parameters; for instance, we might choose to assume the correlation of the responses to be equal between different units. Moreover, in order to perform an overall analysis with post-hoc validity, we might choose a closed testing approach (Marcus et al., 1976), which requires to perform all the $2^{m}$ possible tests, which can be very demanding for growing $m$.
A fast alternative, which is more appealing for large values of $m$, consists of doing marginal standardization of the test statistic. We will start from the effective scores, showing that we are able to derive a valid procedure for both the flipscores approaches. Finally, we will see that we still have a remarkable result for the normal linear model.

### v2

@@ Abstract
In many applied sciences a popular analysis strategy for high-dimensional data is to fit many multivariate generalized linear models in parallel. This paper presents a novel approach to address the resulting multiple testing problem by combining a recently developed sign-flip test with permutation-based multiple-testing procedures. Our method builds upon the univariate standardized flip-scores test which offers robustness against misspecified variances in generalized linear models, a crucial feature in high-dimensional settings where comprehensive model validation is particularly challenging. We extend this approach to the multivariate setting, enabling adaptation to unknown response correlation structures. This approach yields relevant power improvements over conventional multiple testing methods when correlation is present.
@@ 1 Introduction
In high-dimensional data, such as neuroimaging and transcriptomics, it is common to fit many generalized linear regression models in parallel, each with a relatively small sample size (Schaarschmidt et al., 2022; Love et al., 2014; Winkler et al., 2014; Davenport et al., 2022). Usually, the goal of the analysis is to perform hypothesis testing to find relevant associations, typically after adjusting the $p$-values for multiple testing.
@@
This approach comes with several challenges. In the first place, small sample sizes can make classical tests unreliable, especially in nonlinear models where exact methods are not available and tests rely on asymptotic arguments. In such cases, the usual normal approximation of the test statistic can be quite unreliable (Schaarschmidt et al., 2022), especially in the extreme tails that are relevant in the context of multiple testing. Secondly, the generalized linear model makes some crucial assumptions which are difficult to check, especially for nonlinear models with small sample size. In particular, the detection of overdispersion among the assumed models can be quite problematic, a problem which is often exaccerbated as the variance is not constant among the observations. Proper model checking in high-dimensional data is further hampered by the sheer number of models that are fit. Finally, the test statistics of the parallel models are often correlated, due to correlations in the underlying biological measurements. Classical multiple testing corrections (such as Bonferroni-Holm) are designed in order to protect against any correlation structure, but can be very conservative in the presence of such correlations (Gao et al., 2010; Goeman and Solari, 2014; Saffari et al., 2018). Instead, as we shall see, taking the correlation into account allows for increases in power.
Earlier work in this direction have been addressed by Hothorn et al. (2008); Pipper et al. (2012); Bretz et al. (2016); Pallmann et al. (2018) which tackle the multiple testing issue – in contexts that comprise GLMs – by estimating the correlation structure between the test stiatistics and the inference is based on a parametric approach. Unfortunately, they do not allow to deal cases of very high dimensionality as the ones we are tacking in this work.
Some solution is offered by resampling-based methods in the restricted field of linear models. Winkler et al. (2014) provides a thoughtful review of these methods. The asymptotic control of the type I error is good in most of the practical scenarios; but they are not robust to heteroscedastic errors (i.e. misspecified variances) (Anderson, 2001; McArdle and Anderson, 2001) or they are robust in practice, but a formal proof is not provided. The broader field of generalized linear regression models with high dimensional responses offer even less satisfactory tools. As an example, Heller et al. (2008) propose a solution where the effect of the confounders are accounted by matching units with similar euclidean distance in the space of the confounders. Despite being an effective approach, the solution is ad hoc since the quality of the control of the false positives depends on the correct choice of parameters.
@@
In this work we propose a permutation-based multiple testing procedure in combination with the sign-flip score test (Hemerik et al., 2020; De Santis et al., 2022). The test shows reliable behavior for small sample sizes, and has been shown to be robust against general variance misspecification under minimal assumptions. We start by defining a test for a global null hypothesis about a multivariate regression parameter, which guarantees weak control of the familywise error rate (FWER). We extend this to allow for additional inference by embedding the global test in a closed testing procedure (Marcus et al., 1976). This allows us to compute adjusted $p$-values for each individual hypothesis through a multiple testing procedure based on the max-$T$ method of Westfall and Young (1993), which is a dramatic shortcut of the closed testing procedure. In particular this approach guarantees strong control of the FWER, that is, it produces valid adjusted $p$-values for all possible subsets of hypotheses.
Crucially our approaches provides inference by jointly sign-flipping the score contributions allowing the method to adapt to the unknown correlation structure. This is especially useful when strong correlation between the individual test statistics is present; providing a relevant gain in power over alternative methods, as we will show in a simulation.
@@
The paper is organized as follows: Section 2 revisits the sign-flip score test for univariate testing; Section 3 contains the novel contribution of the paper, that is, the multivariate testing extension. Section 4 contains a simulation study; Section 5 contains the conclusions.
@@
Code to implement our methods is available in the flipscores R package (Finos et al., 2022), available from CRAN. Python and Matlab implementations are also available (Davenport, 2023, 2024). Code to reproduce the results of this paper is available at github.com/livioivil/jointest.
@@ 2 One dependent variable
In this section we review the derivation of the univariate sign-flip score test as detailed in Hemerik et al. (2020) and De Santis et al. (2022) before we extend to the multi-variable setting in Section 3.
Given $n$ specify a number of units, which could e.g. correspond to the number of subjects in a given analysis. Then for every $1\leq i\leq n$, let $Y_{i}$ be the target variable which is assumed to belong to the exponential dispersion family, i.e., with a density of the form (Agresti, 2015)
$f(y_{i};\theta_{i},\phi_{i})=\exp\left\{\frac{y_{i}\theta_{i}-b(\theta_{i})}{a(\phi_{i})}+c(y_{i},\phi_{i})\right\},$
@@
where $g(\cdot)$ is the link function, the covariate $X$ is an $n\times 1$ matrix, i.e., a column vector, such that $\beta\in\mathbb{R}$. Further, the nuisance covariates $Z$ are a $n\times(k-1)$ matrix. Let $\eta=X\beta+Z\gamma$ define the linear predictor.
Throughout this section we are interested in testing the null hypothesis $H_{0}:\beta=\beta_{0}$ against a one or two-sided alternative. The other parameters $\gamma$ (and $\phi_{i})$ are thus considered as nuisance parameters. Note that in practice in order to fit generalized linear models it is common to assume that $\phi_{i}=1$ or $\phi_{i}=\phi$, a requirement needed to be able to estimate them (Agresti, 2015). This turns out to an assumption about the variance structure that can be quite restrictive. We do not wish to have this kind of assumption, and so instead rely on score based sign-flipping tests which we introduced in Hemerik et al. (2020) and De Santis et al. (2022) which have robustness against variance misspecification. To do so let
@@
The diagonal matrix of the GLM weights is then defined as $W=DV^{-1}D$ with entries $w_{i}=({\partial\mu_{i}}/{\partial\eta_{i}})^{2}/var(Y_{i})$ (Agresti, 2015). Finally,
$P={W}^{1/2}Z(Z^{\prime}{W}Z)^{-1}Z^{\prime}{W}^{1/2}$
represents the projection matrix for GLMs. Note that $D,V$ and $P$ are functions of the regression parameters $\beta,\gamma$, and we evaluate them under the null hypothesis $H_{0}$ i.e. with $\beta=\beta_{0}$ and $\gamma=\gamma_{0}$.
We will consider a test based on the effective score which, when the estimated nuisance parameter $\hat{\gamma}$ is plugged in, is defined as
$n^{1/2}S^{*}_{\hat{\gamma}}=X^{\prime}{W}^{1/2}(I-{P}){V}^{-1/2}(Y-\hat{\mu})=\sum_{i=1}^{n}\nu^{*}_{\hat{\gamma},i}.$
The effective score can be interpreted as the residual from the projection of the marginal score for $\beta$ onto the space spanned by the nuisance scores (Marohn, 2002). Importantly the effective score can be written as a sum of $n$ elements $\nu^{*}_{\hat{\gamma},i}$, which we call the score contributions.
In what follows we shall adopt the same model assumptions and estimation strategy as in De Santis et al. (2022). In particular, the link function is crucially assumed to be correctly specified which implies that we are able to consistently estimate the regression parameter $\gamma$.
The test is performed by means of random sign flipping of the score contributions. Define $g_{1}=(1,\dots,1)\in\mathbb{R}^{n}$ and let $g_{2},...,g_{w}$, for some number of flips $w\in\mathbb{N}$, be sampled with replacement from $\{-1,1\}^{n}$, unless stated otherwise. The sign-flipped effective score statistics are then defined as
$S^{*j}_{\hat{\bm{\gamma}}}$ $=n^{-1/2}\sum_{i=1}^{n}g_{ji}\nu^{*}_{\hat{\bm{\gamma}},i},$
where the superscript $j$ denotes that the $j$-th flip $g_{j}$ has been applied.
For conciseness it is helpful to write the sign-flipped effective score statistic in this matrix notation. To do so note that the effective score is the product of an $1\times n$ vector $X^{\prime}{W}^{1/2}(I-{P}){V}^{-1/2}$ and an $n\times 1$ vector $(Y-\hat{\mu})$. For each $1\leq j\leq w$, letting $G_{j}$ be the diagonal $n\times n$ matrix with diagonal entries $g_{j1},\dots,g_{jn}$, the sign-flipped test statistic can thus be written as
$S^{*,j}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}{W}^{1/2}(I-{P}){V}^{-1/2}G_{j}(Y-\hat{\mu}).$
@@
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}=S^{*j}_{\hat{\bm{\gamma}}}$ and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted test-statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{0}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
We will now recall the definition of the test proposed in De Santis et al. (2022), which is a recent upgrade of the test above. This recent adaptation typically improves the small-sample performance of the test by standardizing the effective score. We call the resulting test statistic the standardized score statistic, defined as
$S_{\hat{\bm{\gamma}}}=S^{*}_{\hat{\bm{\gamma}}}/Var\{S^{*}_{\hat{\bm{\gamma}}}\}^{1/2},$ (1)
@@
$Var\{S^{*}_{\hat{\bm{\gamma}}}\}=n^{-1}X^{T}W^{1/2}(I-P)W^{1/2}X.$
The corresponding sign-flipped test statistic, for a generic flip matrix $G_{j}$, is defined as
$S^{j}_{\hat{\bm{\gamma}}}=S^{*,j}_{\hat{\bm{\gamma}}}/Var\{S^{*,j}_{\hat{\bm{\gamma}}}\}^{1/2},$ (2)
@@
$Var\{S^{*,j}_{\hat{\bm{\gamma}}}\}=n^{-1}X^{T}W^{1/2}(I-P)G_{j}(I-P)G_{j}(I-P)W^{1/2}X.$
In practice, this variance has to be estimated, replacing $W$ with an estimate $\hat{W}=\hat{D}\hat{V}^{-1}\hat{D}$, where $\hat{D},\hat{V}$ are estimates of $D,V$. This introduces an asymptotic approximation of lower order with respect to the parameter estimation. Indeed, in a first-order analysis the use of the true or the estimated variance is equivalent (Pace and Salvan, 1997). All results in the rest of the paper will be formulated in terms of $W$, but the results will hold up to the first order in terms of $\hat{W}$ (De Santis et al., 2022).
@@
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}=S^{j}_{\hat{\bm{\gamma}}}$ and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{0}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
Some robustness properties of this test are discussed in De Santis et al. (2022). In particular, the proposed test is proven to be asymptotically exact in case of general variance misspecification under minimal assumptions.
3 Multiple dependent variables
We now generalize the results of Section 2 to the case of multiple generalized linear models fitted in parallel. Suppose there are $m\geq 2$ dependent variables $Y^{1},\dots,Y^{m}$ and that for each response $Y^{l}$ we have $n$ independent observations $Y^{l}_{1},\dots,Y^{l}_{n}$, which follow some model in the exponential dispersion family. We consider $m$ null hypotheses $H_{1},\dots,H_{m}$, where $H_{l}$ is the hypothesis that $\beta^{l}=\beta^{l}_{0}$. For the $l$-th model we assume the same assumptions of Section 2.
@@ 3 Multiple dependent variables
$S^{*,l}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}({W}^{l})^{1/2}(I-{P}^{l})({V}^{l})^{-1/2}(Y^{l}-\hat{\mu}^{l})=n^{-1/2}\sum_{i=1}^{n}\nu^{*l}_{\hat{\bm{\gamma}},i},$
where the superscript $l$ indicates that the corresponding quantities have been for the $l$th model as in Section 2. Analogously to Section 2, we define the sign-flipped effective score test statistic for a generic flip matrix $G_{j}$ as
$S^{*,j,l}_{\hat{\bm{\gamma}}}=n^{-1/2}X^{\prime}({W}^{l})^{1/2}(I-{P}^{l})({V}^{l})^{-1/2}G_{j}(Y^{l}-\hat{\mu}^{l})=n^{-1/2}\sum_{i=1}^{n}g_{ji}\nu^{*l}_{\hat{\bm{\gamma}},i}.$
Using the same sign-flip for each model allows the procedure to preserve the dependence structure. Moreover, we have
$Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}=X^{\prime}({W}^{l})^{1/2}(I-{P^{l}})G_{j}(I-{P^{l}})G_{j}(I-{P^{l}})({W}^{l})^{1/2}X.$
This allows us to define the standardized score as
$S^{j,l}_{\hat{\bm{\gamma}}}=S^{*,j,l}_{\hat{\bm{\gamma}}}Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}^{-1/2}.$
@@
$H_{L}=\bigcap_{l\in L}H_{l}:\beta^{l}=\beta^{l}_{0}$ (3)
where $L$ is a fixed set containing any combination of the $m$ hypotheses. Note that $H_{L}$ can be either a global or a partial null hypothesis for a subset of parameters. For simplicity, we will derive the procedure for the global null hypothesis, but it is exactly analogous for partial null hypotheses of any kind.
A first idea to perform the test might be given by the following full standardization approach which builds a global test statistic based on the form of a Mahalanobis distance as follows. Let $S^{*,j,L}_{\hat{\bm{\gamma}}}$ be the $m$-dimensional vector with elements $S^{*,j,l}_{\hat{\bm{\gamma}}}$, $1\leq l\leq m$. Let $Var\{S^{*,j,L}_{\hat{\bm{\gamma}}}\}$ denote the corresponding variance-covariance matrix, which has dimension $m\times m$ and is assumed (for the moment) known. Let the joint test statistic be
$T_{j,L}^{n}=\left(S^{*,j,L}_{\hat{\bm{\gamma}}}\right)^{\prime}Var\left\{S^{*,j,L}_{\hat{\bm{\gamma}}}\right\}^{-1}\left(S^{*,j,L}_{\hat{\bm{\gamma}}}\right).$ (4)
for $1\leq j\leq w$. It can be proven that an asymptotic $\alpha$-level test can be built with a reasoning analogous to Theorem 2.
However there are some issues with this approach. In particular for each flip it is necessary to invert an $m\times m$ matrix, which becomes infeasible for large values of $m$. Moreover, it requires estimating the correlation between responses, which must be assumed to be known except for a limited number of parameters; for instance, we might choose to assume the correlation of the responses to be equal between different observations. Finally, in order to perform an overall analysis with post-hoc validity, we might choose a closed testing approach (Marcus et al., 1976), which requires to perform all the $2^{m}$ possible tests, which can be very demanding for growing $m$.
A fast alternative, which is more appealing for large values of $m$, consists of doing marginal standardization of the test statistic. We will start from the effective scores and then proceed to the standardized scores showing that we are able to obtain asymptotically valid inference for both procedures.